OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1 = 12 and a13 = –60?

OpenStudy (anonymous):

tn= a+(n-1)d -60= 12+(13-1)d -60=12+12d -72=12d d=-6 tn=a+(n-1) = 12+ (32-1)-6 =12+ (31x-6) = 12-186 =-174

OpenStudy (anonymous):

thanks so much!

OpenStudy (anonymous):

no problem